A metre long tube open at one end, with a moveable piston at the other end, shows resonance with a fixed frequency source (a tunning fork of frquency 340 Hz) when the tube length is 25.5 cm or 79.3 cm. Find the speed of sound in air at the temperature of the experiment. (Ignore end correction)

Grade:12

1 Answers

Vikas TU
3 years ago
when successive resonance frequencies are given we consider the twice of the length difference at which these occur to be the wavelength of the waves .
Here, length of first resonance occurs ( L1) = 25.5 cm =
Length of 2nd resonance occurs ( L2) = 79.3 cm
So, wavelength = 2 × ( L2 - L1)
= 2 × ( 79.3 - 25.5)
= 2 × 53.8
= 107.6 cm
= 1.076 m
Now, speed of sound = frequency × wavelength
= 340 × 1.076 m/s
= 365.84 m/s